An aqueous solution of Borax (Na2B4O7) is sometimes used to fireproof wood. 


What is the molarity of 2.50 L of solution that contains 1.85 mol of Na2B4O7?  


1
Expert's answer
2022-03-30T03:18:04-0400

CM = n / V

CM (Na2B4O7) = 1.85 / 2.50 = 0.74 M